Reports 1-1 of 1 Clear search Modify search
AdV-DAQ (Calibration)
grimaud, rolland - 19:02 Thursday 03 April 2025 (66503) Print this report
Power calibration at NE, measurement of reflection bench mirror loss and measurement of NE Rx sphere noise on 2025-04-03

Plan for the day 

- Test of angle of incidence and mirror loss on reflection bench of NE PCal

- Test of angle of incidence and mirror loss on reflection bench of WE PCal

- Test of NE_Rx at WE PCal

 

Test of the angle of incidence and mirror loss on NE reflection bench mirror

1- mirror loss measurement
Around 7h20UTC we opened the reflection bench and placed the WSV sphere in front of the M3 mirror to measure the beam right after the viewport.
At 7h25 UTC, with laser at 1.3 W  we measured a power on WSV = 1.29882 +- 3.5e-4 W
At 7h36 we put WSV sphere at its initial position and measure a power on WSV =1.2966 +- 3.9e-4 W 

Those measurement are giving a value of mirror loss of 0.17% (comparable to LAPP measurement of 0.12% and last measurement made on site in June 2023 of 0.12% also)
This seem to indicate that most of the discrepency between NE and WE are coming from the viewport rather than the M3 mirror.

 

2- Visual inspection of the viewport

In order to visually inspect the viewport on the reflection bench we opened the diaphragm and shined a flash light onto it, the inspection was not conclusive as it seem to be some minimal scratch and dust but not particularly in the middle of the viewport.
After this we tried closing the diaphragm as it was before.

A first measurement made at 7h52UTC gave us a value of power on WSV =  1.29576 +- 3.5e-4 W

We tried closing the diaphragm more at 7h55 UTC and measured a power on WSV slightly lower.
We put the diaphragm back at it initial setting at 7h58UTC. 

 

At 8h05 UTC We closed NE Rx bench, with WSV at Rx spot and the laser at 1.3 W to redo a short calibration measurement for the photodiodes to check if touching the diaphragm changed the calibration.

 

Calibration AFTER playing with the Rx diaphragm. (PCal NE lines stopped)

THe background measurement is done using the same data than for the first calibration : 1427652018 for 3600s

The time period for this calibration measurement is from 10h10UTC to 13h39 : 1427710218 for 12540s

 

old gain

new gain diff old offset new offset
Tx_PD1 -0.775251 -0.772392 0.37% -0.000757 -0.000768
Tx_PD2

-0.858945

-0.858746 0.023% -0.000843 -0.000861

There is no significant change from the first calibration performed on the 2025-04-03 before touching the diaphragm. The plot of the voltages ratio from WSV to the photodiodes is in plot1.

 

Test of angle of incidence and mirror loss on reflection bench of WE PCal

We arrived at WE building at 8h25, and stoped the laser and Rx power supplu at around the same time.

1- Test of Rx blue box value when not plugged but still powered

8h30 : 15 V power supply stopped (blue box not stopped)
~8h33 :  Rx cable unplugged on Rx sphere
8h33 : blue box stopped, power supply switched on  → WE_Rx_PD1_DC goes up to 3.6 W constant . ADC saturation?
8h36: blue box switched on, still constant value.
8h37 : blue box and power supply off . Value back to ~0
8h40 : Rx sphere plugged, then power supply on, and blue box on-> value ~0.

The same effect was seen later with WSV when retrieving it from NE, in plot 2 we can see three steps of power value, first when the laser was on at 1.3W reading around -3.5 V.
Then when the laser was turned off and the WSV blue box turned off with a value around 0V and finally when the cable was disconnected from the sphere where we have a high reading of around 3 V (see plot 3) similar to the Rx case seen before.


2- Measurement of mirror loss
During this test we noticed some glitches in the data, more investigation will be done but not mirror loss measurement were performed

Measurement of Rx_NE noise at WE

8h43 : laser back to 1.3 W.  

8h45 : Stop PCal lines on WE.

10h45 : start of the noise level measurement of Rx_WE
10h47 : stoped laser and stop blue box of Rx_WE
10h50 : replace the Rx sphere of WE by the Rx sphere of NE to test the noise of Rx_NE.
10h58 -> 11h05 : start blue box, start laser and ctrl loop at 1.3 W. -> Channel PCAL_WE_Rx_PD1_DC is now reading the power using the Rx sphere from NE.
11h10 : WE_Rx sphere put back, laser switched on to 1.3 W
11h16 : put back calibration lines on NE and WE

In plot4 we can see the comparison of the noise level of the Rx_WE and Rx_NE sphere with the photodiodes as witness channels.
We see that the increased noise level for Rx_NE is also seen when reading it with Rx_WE electronics indicating that the higher noise is coming directly from the NE sphere itself.

Retrieving WSV from NE

Arrived at the NE building around 13h30 UTC, turned off laser around 13h35UTC

FInished removing WSV and its electronic at 14h05 UTC

Turned Rx and laser back on at 14h10 UTC

Images attached to this report
Comments to this report:
rolland, grimaud, boschi, razzano et al. - 7:19 Friday 04 April 2025 (66508) Print this report

In order to perform the optical losses measurements at WE around 8h-10h UTC, we have stopped the WE PCal permanent lines. During these measurements, we have seen glitches on the power sensors.

We initially though it may be related to the SAT activities to install a PDU on a rack close to the calibration rack (see logbook 65501), and in particular related to switching on the motor drivers plugged on this PDU. However, the conclusions are not clear finally.

Here are some plots as a reminder:

  • figure 1: a short period with glitches during ~3 seconds. The peak-to-peak values of the power during the  period without noise are 1.299 to 1.301 W , it increases up to 1.295 to 1.305 W in the noise period.
  • figure2: zoom on the same period.
  • figure 3: another period, of about 1.5 minutes, with glitch rate increasing, and then they suddenly stopped.
  • figure 4: around 9h14 UTC, the glitches seemed to appear at the time the SAT team started the motors. This very glitchy period lasted all along the motor tests, but did not stop at the end. 
  • Around 9h48-9h54, we restarted the PCal processes and the glitches were not there after restart.
  • Figure 5: At 9h56, we made a test: SAT team restarted the motor drivers, and glitches re-appeared with increasing rate. They were stop 1 or 2 minutes later, and the glitch rate decreased and suddenly they stopped.... however, glitches restarted a bit later, without particular action from SAT.
  • FIgure 6: summary of the morning activities at WE from the trend of the channels, where the presence or absence of glitches is well visible looking at the min/max values.
  • SAT team left WE building around 10h20-10h30.
  • 10h30-10h40: stopped all PCal hardware (switch off the Hameg power supply and the DaqBox, ), and restarted all, to compare the noise of the WE and of the NE sphere until 11h30 (see previous logbook entry).
  • Figure 7 : at 11h40, put back the permanent lines. The amplitude of the lines is much larger than the glitches (the bottom plot has been zoomed on y-axis to see some glitches before the calibration lines were restarted).

We did not notice such glitches during the run. However, they may have been present, but not visible since the calibration lines were always on.

In the data from this morning, it looks like the glitch rate was higher and over longer periods when there were activities around the calibration and SAT motor racks in the WE building, but it is not very conclusive.

The PCal power loop is using the single-ended output of the fastDAC. It is known that (at least some) fastDAC channels have some oscillation at ~MHz high frequency.  A past quick test at NE (logbook 66132) to use differential output was not useful about the noise level. But could it generate these small glitches?

To check if the situation is similar for NE PCal, it will be worth stopping the NE PCal lines also for a long period during the WE mirror replacement next week.

 

Images attached to this comment
Search Help
×

Warning

×